OpenStudy (anonymous):

Write the slope-intercept form of the equation of the line which passes through the points (1, 6) and (2, 9). (

OpenStudy (texaschic101):

first you need to find the slope by using the slope formula.. slope(m) = (y2 - y1) / (x2 - x1) (1,6) x1 = 1 and y1 = 6 (2,9) x2 = 2 and y2 = 9 now we sub slope(m) = (9 - 6) / (2 - 1) slope(m) = 3/1 or just 3 now we use y = mx + b (we already have all the numbers except for b, the y intercept y = mx + b slope(m) = 3 you can use either set of points, it does not matter. I will use (1,6) x = 1 and y = 6 now we sub and find b 6 = 3(1) + b 6 = 3 + b 6 - 3 = b 3 = b so our equation of the line is : y = 3x + 3 any questions ?

OpenStudy (texaschic101):

good to hear....just to let you know, instead of using y = mx + b, it could have been solved with the point-slope formula, y - y1 = m(x - x1) and then turned to slope intercept form. I just find it easier to use y = mx + b unless it specifically asks for point-slope form